A population of values has a normal distribution with μ=152.3 and σ=54.2. You intend to draw a random sample of size n=245.

Find the probability that a single randomly selected value is between 141.2 and 145.4.
P(141.2 < X < 145.4) =

Find the probability that a sample of size n=245 is randomly selected with a mean between 141.2 and 145.4.
P(141.2 < M < 145.4) =

Enter your answers as numbers accurate to 4 decimal places. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Solution :

Given that ,

mean = = 152.3

standard deviation = = 54.2

P(141.2 < x < 145.4) = P[(141.2 - 152.3)/ 54.2) < (x - ) /  < (145.4 - 152.3) / 54.2) ]

= P(-0.205 < z < -0.127)

= P(z < -0.127) - P(z < -0.205)

= 0.4495 - 0.4188

= 0.0307

P(141.2 < x < 145.4) = 0.0307

n = 245

M = 152.3

M = / n = 54.2 / 245 = 3.4627

P(141.2 < M < 145.4) = P((141.2 - 152.3) /3.4627 < (M - M) / M < (145.4 - 152.3) / 3.4627)

= P(-3.206 < Z < -1.993)

= P(Z < -1.993) - P(Z < -3.206)

= 0.0231 - 0.0007

=0.0224

P(141.2 < M < 145.4) = 0.0224